OpenStudy (shamil98):

Help with differentiation please.. using the chain rule. y = (3x+1)^2

OpenStudy (shamil98):

Would y = 2(3x+1)?

hartnn (hartnn):

derivative of x^2 will be 2x but using chain rule, derivative of (3x+1)^2 will be 2(3x+1) * d/dx(3x+1) = ... ?

OpenStudy (shamil98):

= 2(3x+1) (3) right?

hartnn (hartnn):

yes :)

OpenStudy (shamil98):

yay! thanks hartnn :)

hartnn (hartnn):

welcome ^_^

OpenStudy (timaashorty):

DANG I MISSED THIS, OR I COULD OF HELPED YOU TOO (: